# Kiet A. NguyenAdheesh JuvekarTianjiao YuMuntasir WahedIsmini Lourentzou

### v1

@@ C.2 Evaluation Metrics
To evaluate the performance of models in predicting semantic labels for the segmented objects and parts, we compute Semantic Similarity and Semantic IoU. Semantic Similarity (SS) is the semantic similarity between predicted and true labels in the SentenceBERT [48] embedding space, following past works [67, 7], while Semantic IoU (S-IoU) computes the word overlap between predicted and ground truth labels, i.e.,
[…]
where $V(y)$ represents the set of words comprising label $y$.
Appendix D PartImageNet Modifications
While PartImageNet provides high-quality segmentation masks across a diverse range of object classes, its categories are often too abstract and not commonly used in natural language due to their broad scope. For instance, PartImageNet classifies all four-legged animals under the “quadruped” supercategory. Although technically accurate, such classifications are too generic for practical use in everyday language. To address this, we manually selected the most commonly used object class associated with each category provided by ImageNet. The dataset includes WordNet synset IDs (e.g., “n02071294”) that correspond to specific object classes. Using these synset IDs, we extracted the hierarchical path of each class within the WordNet graph. For example, the synset ID “n02071294” maps to the following WordNet path: living_thing $\to$ organism $\to$ animal $\to$ chordate $\to$ vertebrate $\to$ mammal $\to$ placental $\to$ aquatic_mammal $\to$ cetacean $\to$ whale $\to$ toothed_whale $\to$ dolphin $\to$ killer_whale. From this path, we selected whale as the representative object category for this class. We repeat this process for all object classes provided by PartImageNet. The object classes we use alongside the original supercategory in parentheses are shown in Table 6.
Appendix E Additional Experiments
E.1 Computational Efficiency Evaluation
For multi-image tasks, it is essential that the model is efficient so that inference and training do not consume more time as the number of images increases. Therefore, we opt to use Q-Former to query image embeddings, resulting in only 32 tokens per image, which is 8 times fewer tokens than LISA (256 tokens) and 18 times fewer tokens than GLaMM (576 tokens), reducing TFLOPs over LISA by $42.84\%$ and over GLaMM by $22.62\%$ (LISA:19.63 TFLOPs vs. GLaMM:14.5 TFLOPs vs. Calico:11.22 TFLOPs). This corresponds to an $\sim$1.75$\times$×speed-up and $19.39\%$ relative performance gain over LISA. These results demonstrate that Calico achieves significant improvements in both computational efficiency and performance gains compared to baselines.
E.2 Per-Task Experimental Results
In Table 5, we report experimental results decomposed into the 3 MixedParts subtasks (common objects, common parts, and unique parts). The results delineate the incremental difficulty of the tasks by decreasing scores across all models, with common objects being the easiest task, followed by common parts, and finally unique parts as the hardest task. Calico demonstrates superior performance across all 3 tasks, with large leaps in scores across all metrics compared to the next best baseline.
E.3 Low-Tail Performance
0-5.05.0-10.010.0-15.015.0-20.020.0-25.00020202020404040406060606080808080(a) Object Frequency (%)Recall
0-5.05.0-10.010.0-15.015.0-20.020.0-25.025.0-30.000101010102020202030303030404040405050505060606060(b) Common Part Frequency (%)Recall
0-5.05.0-10.010.0-15.015.0-20.020.0-25.025.0-30.00010101010202020203030303040404040(c) Unique Part Frequency (%)
Figure 7: Low-tail analysis of the three subtasks in MixedParts. (a), (b), and (c) demonstrate Calico’s performance across varying object and part frequency levels, showing Calico remains robust under different object-part frequencies.
We show results for low-tail analysis of Calico’s performance in Figure 7. We compute the model’s average recall across different object-part frequency levels on the three MixedParts subtasks. We observe no clear correlation between object-part frequencies and model performance. This highlights Calico’s robust performance across diverse object-part classes in the MixedParts image sets.
Figure 8: MixedParts: Object Instance Distribution.
Figure 9: MixedParts: Part Instance Distribution.
Figure 10: MixedParts: Common Parts Instance Distribution.
Figure 11: MixedParts: Unique Parts Instance Distribution.
@@ E.3 Low-Tail Performance
Table 6: Object and part taxonomies by dataset (continued in next page).

### v2

@@ D.2 Per-Task Experimental Results
Sparkles: Extracting all objects from both images: What are the objects in IMAGE#1<Img><ImageHere></Img> and IMAGE#2<Img><ImageHere></Img>? Reply with only the objects that visible in the images. Only talk about visible features, and limit output to the different objects. Make sure that each object is covered and described properly. Please talk about both the images and do not repeat objects. Extracting all parts from both images: What are the parts of objects in IMAGE#1<Img><ImageHere></Img> and IMAGE#2<Img><ImageHere></Img>? Reply with only the parts of objects that visible in the images. Only talk about visible features, and limit output to the different object and parts. Make sure that each part is covered and described properly. Please talk about both the images and do not repeat parts. GPT4o: Common objects: You will be given description of two images. Understand the description of the images, make a list of the objects in both the images, and identify common objects. If you do not find any common objects, you can leave it empty. Return these as a JSON object, strictly in the following format, output nothing else: { common_objects’: [’object1’, object2’, ...] } replacing objectX with the actual object names. Or if there are no common objects: { common_objects’: [] } Heres the description of both the images: <Sparkles’ outputs>. Please maintain the JSON output format with object names for both images. Common parts: You will be given description of two images. Understand the description of the objects, make a list of the objects and their parts in both the images, and identify the common parts from all the objects in the images. If you do not find any common parts, you can leave it empty. Return these as a JSON object, strictly in the following format, output nothing else: { common_parts’: [ {’image1’: [object1, part1], image2’: [object1, part1]}, {’image1’: [object1, part2], image2’: [object1, part2]}, ... {’image1’: [object2, part1], image2’: [object2, part1]}, ...] } replacing objectX and partX with the actual object and part names. Or if there are no common parts: { common_parts’: [] } Heres the description of both the images: <Sparkles’ outputs>. Please maintain the JSON output format with both object and part names for both images. Unique parts: You will be given description of two images. Understand the description of the objects, make a list of the objects and their parts in both the images, and identify the unique parts from all the objects in the image. If you do not find any unique parts, you can leave it empty. Return these as a JSON object, strictly in the following format, output nothing else: { unique_parts_image1’: [ [’object1’,’part1’], [’object1’,’part2’], [’object2’,’part1’], ... ] unique_parts_image2’: [ [’object1’,’part1’], [’object1’,’part2’], [’object2’,’part1’], ... ] } replacing objectX and partX with the actual object and part names. Or if there are no unique parts: { unique_parts_image1’: [], unique_parts_image2’: [] } Heres the description of both the images: <Sparkles’ outputs>. Please maintain the JSON output format with both object and part name for both images. LISA: Repeat for all common objects: Can you please segment <object> in this image? Repeat for all (common or unique) parts: Can you please segment <part> of <object> in this image?
Figure 10: Prompts for the Cascaded Pipeline.
Figure 11: Top-30 Most Frequent Object Categories in MixedParts.
Figure 12: Top-30 Most Frequent Part Categories in MixedParts.
Figure 13: Top-30 Most Frequent Common Parts in MixedParts.
Figure 14: Top-30 Most Frequent Unique Parts in MixedParts.
@@
Table 6: Object and Part Taxonomies by Dataset (to be continued on next page).
@@
Appendix E Limitations
This research introduces Calico, a model designed for part-focused semantic co-segmentation, incorporating several novel features. While our contributions include the proposal of a valuable new task and the release of a supporting dataset to encourage future research, there are a few limitations and assumptions that warrant discussion. First, Calico assumes that semantically meaningful correspondences can be established between similar object parts across different images based on visual features. However, this assumption may break down in scenarios where visually similar parts serve different functions. Like many machine learning models, Calico demonstrates promising results on curated datasets, but its generalizability to complex, real-world environments remains to be fully validated. Future work should address these limitations, potentially through adaptive learning strategies that improve robustness and scalability across diverse, real-world applications.
Appendix F Broader Impact
This work introduces several significant advancements with broad societal implications—both positive and cautionary. The integration of object and part comparison methods holds promise for numerous applications in areas such as automated quality control in manufacturing, enhanced image-based search engines, and more sophisticated systems for digital content management and creation. For example, the proposed new task and models can automate and improve the accuracy of tasks that require detailed visual comparisons, such as quality assurance in manufacturing, potentially increasing efficiency while reducing human error. Calico offers a more fine-grained understanding of images by identifying and segmenting objects across images based on their constituent common or unique parts. This can greatly aid in fields such as robotics, where agents require a detailed understanding of object parts for precise manipulation or obstacle avoidance tasks. Furthermore, in medical imaging, fine-grained visual decomposition can assist in detailed diagnostic tasks. Improved part-level understanding can also improve accessibility technologies, such as software for the visually impaired, by providing more descriptive and accurate visual summaries.
However, reliance on visual segmentation alone may introduce risks in tasks where non-visual attributes are crucial. Objects with similar appearance but different material properties (e.g., plastic vs. metal) may be misinterpreted, potentially affecting tasks in high-stakes settings such as surgery or industrial handling of fragile materials. Depending on the application, the integration of other sensory data inputs (tactile, thermal, or acoustic sensors), together with visual data, would be beneficial in mitigating this risk. Future work can also design new datasets and models that consider metadata about material properties and other attributes, and ensure that the model is trained not just to recognize visual object and part similarities/differences, but also to associate these features with the correct properties. Finally, human-in-the-loop solutions that incorporate human oversight can ensure critical decisions are validated, improving reliability and trustworthiness in deployment.
Appendix G Image Attributions
The three images in Figure 7, used for qualitative illustration of our model’s in-context capabilities, are credited respectively to Erik Mclean, Pixabay, and Trace Constant on Pexels. All other images shown in this paper are sourced from publicly available datasets.